Three cards are drawn from a deck of 52 cards without replacement. Find the probability that all are of the same color
1
Expert's answer
2020-03-02T15:07:54-0500

In a deck there are a 26 red cards and 26 black card. We need the probability that both are red or both are black.

P(RRR) or P(BBB)

P(RRR) =26/52*25/51*24/50

=2/17

P(BBB)=26/52*25/51*24/50

=2/17

The probability that all are of the same color=2/17+2/17

=4/17